# Lifelong and Universal Machine Learning Potentials for Chemical Reaction Network Explorations

### v1

@@ 2.2 Lifelong Adaptive Data Selection
If the exclusion strike counter $X^{r}$ reaches its limit $N_{\mathbf{X}}$, $S_{\mathrm{hist}}^{r,E/F}$ is set to $-3$ and the structure is excluded due to very large errors. If the upper threshold for $S_{\mathrm{hist}}^{r}$ as defined in Algorithm 2 is exceeded, $-2$ is assigned to $S_{\mathrm{hist}}^{r,E/F}$ to exclude the structure due to steadily large errors for many training steps. Both assignments mean that the respective training data point is inconsistent with the majority of data. If the lower threshold for $S_{\mathrm{hist}}^{r}$ is exceeded, $S_{\mathrm{hist}}^{r,E/F}$ is set to $-1$, classifying a redundant structure, i.e., a structure that has been well represented for many steps. To avoid that too many structures are classified as redundant in the same training step, a maximal fraction of the new redundant training structure per step $p_{\mathrm{redun}}^{\mathrm{max}}$ can be set in addition to Algorithm 2. In this way, only
@@
randomly selected structures of all structures exceeding the lower threshold are assigned to be redundant and $S_{\mathrm{hist}}^{r,E/F}$ of the remaining structures is divided by $F_{--}$.
To update the fraction of good data $p_{\mathrm{good}}$, the new loss contributions first replace the respective old ones in $\mathbf{L}_{\mathrm{old}}^{E/F}$. Subsequently, the mean $\overline{L}_{\mathrm{new}}^{E/F}$ of this vector is calculated for all available training data and those assigned to be redundant. If the new means of energies and forces are not larger than the respective old values, $p_{\mathrm{good}}$ will decrease by $p_{\mathrm{good}}^{\mathrm{max}}\,N_{p}^{-1}$. Otherwise, it increases by the same value. The resulting $p_{\mathrm{good}}$ has a lower bound of 00 and an upper bound of $p_{\mathrm{good}}^{\mathrm{max}}$. Consequently, $p_{\mathrm{good}}$ can have $N_{p}+1$ different values. Finally, $\overline{L}_{\mathrm{old}}^{E/F}$ is overwritten by $\overline{L}_{\mathrm{new}}^{E/F}$.
@@
For an efficient integration of additional data at a training stage where some structures are already well represented, Algorithm 3 can replace lines 3 to 6 of Algorithm 2. The main idea behind this data integration is that the additional data can have the maximum training probability for several training steps, while they cannot be excluded during these steps. In this way, the lMLP can learn the new structures fast, even if the representation quality is very different between the previous and new training data. Without this integration algorithm, the risk of exclusion is high for new data because the typically low errors for the majority of old data lead to a fast increase in the exclusion strike counters and the adaptive selection factors of new data with high errors. In addition, the new data initially do not affect the update of the selection-determining properties of the old data to obtain a stable assessment of the representation quality.
@@
$\quad\quad\overline{L}_{Ti}^{E/F}\leftarrow\mathrm{mean}\left(\mathbf{L}_{Ti}^{\backslash\mathrm{NaN},S_{\mathrm{hist}}^{r,E}\geq-1,E/F}\right)$
@@
[…]
@@
[…]
@@
The following part of Algorithm 4 has to be repeated for each structure $r^{\prime}$ that has been identified in the current step to be inconsistent. Initially, the non-normalized probabilities $\mathbf{P}^{r^{\prime}}$ are calculated using the current $P_{\mathrm{loss}}^{r^{\prime}}$ and $p_{\mathrm{good}}$ and $\mathbf{F}_{\mathbf{S}}$ or $\mathbf{F}_{\mathbf{X}}$. The latter depends on which exclusion reason applies for $r^{\prime}$. Afterwards, the average training frequency is calculated for each entry in $\mathbf{P}^{r^{\prime}}$, whereby a structure can only be trained once in a training step. The cumulative sum minus the value of the first vector entry yields the vector of backtracking steps $\mathbf{t}_{\mathrm{BT}}$. The current epoch $n_{\mathrm{epoch}}$ minus $\mathbf{t}_{\mathrm{BT}}$ corresponds to the most probable training steps of structure $r^{\prime}$ under the aforementioned assumptions. For these steps, the decay hyperparameters $\boldsymbol{\upbeta}_{1}$ of the CoRe optimizer are calculated. Subsequently, the backtracking factors $F_{\mathrm{BT}}^{1}$ and $F_{\mathrm{BT}}^{2}$ can be obtained, i.e., the sum of the weights of all loss gradient contributions of structure $r^{\prime}$ in the steps $n_{\mathrm{epoch}}-\mathbf{t}_{\mathrm{BT}}$ can be calculated using the decay hyperparameters $\boldsymbol{\upbeta}_{1}$ and $\beta_{2}$.
@@
For the lADS hyperparameters, we recommend these settings: $p_{\mathrm{fit}}=\tfrac{1}{30}$, $S_{\mathrm{hist}}^{\mathrm{min}}=0.1$, $S_{\mathrm{hist}}^{\mathrm{max}}=100$, $T_{i}=\{0.9,2.5,4.0,7.5\}$, $N_{--}=10$, $N_{-}=30$, $N_{+}=100$, $N_{++}=300$, $N_{\mathbf{X}}=15$, $p_{\mathrm{redun}}^{\mathrm{max}}=0.02$, $p_{\mathrm{good}}^{\mathrm{max}}=\tfrac{2}{3}$, $N_{p}=20$, and $N_{\mathbf{I}}=30$. The data integration algorithm (Algorithm 3) can be applied as soon as new data are added for the first time. Hence, Algorithm 3 needs to be disabled only if training is started from scratch. In addition, the hyperparameter for balancing the energy and force loss contributions was $q=250$ in this work. These settings were applied unless otherwise stated.
2.3 Universal Machine Learning Potentials
The goal of uMLPs is to achieve out-of-the-box applicability for chemically meaningful structures. Then, further retraining is not necessary and expensive reference calculations can be avoided during application. To achieve this goal, models with universal structure representations are trained on very large and diverse reference data sets.
ANI [67] is one of the initial attempts to create a more general MLP. Still, the applicability of ANI-1x [63] and ANI-2x [68] are, among other restrictions, limited to four and seven chemical elements, respectively. Graph representations can remove this limitation, as demonstrated, for example, by M3GNet [44] and MACE [69]. These MLPs were trained on data from the Materials Project [70, 71] and can be universally applied to almost the entire periodic table. In recent years, many more general and universal models have been proposed, such as ANI-1xnr [72], AIMNet2 [73], TeaNet [74], PreFerred Potential [75], TeaNet/PreFerred Potential [76], ALIGNN-FF [77], GNoME [78], MatterSim [79], and CHGNet [80] (in increasing order of covered elements).
Current uMLPs have already shown remarkable success in various benchmarks providing accurate and stable simulation results [46, 81]. However, these benchmarks have also revealed that there are still a lot of applications, for which the accuracy of the uMLPs is not sufficient and further training is required. Fine-tuning of the universal foundation models on specialized data can often be a solution to increase the accuracy to the desired level. However, to reach the final goal of out-of-the-box applicability, the diversity and size of the uMLP reference data sets needs to be increased. We note that the combination of an MLP with a semi-empirical method as in AIQM1 [82] can also increase transferability and accuracy, while computational efficiency is at an affordable level.
Beyond the applicability for different chemical elements, the total charge and spin states have to be handled by the uMLP as well. Current uMLPs are trained on data of neutral systems in the electronic ground state which compromises predictions on charged systems. In addition, for a comprehensive representation of chemical space, training data need to cover different aggregation states and various conformations of the system, including out-of-equilibrium structures. The focus of uMLPs is typically either on gas-phase molecules or on solid materials. Attempts have been made to integrate transitions between equilibrium structures in general training data sets [83].
[…]
@@ 3.1 Exploration of Chemical Reaction Networks
For the CRN exploration we applied our freely available open-source Software for Chemical Interaction Networks (SCINE) [86], especially the SCINE modules Chemoton (version 3.0.0) [23, 87], Molassembler (version 2.0.0) [88, 89], ReaDuct (version 5.0.0) [90, 91], Puffin (version 1.2.0) [92], and Database (version 1.2.0) [93]. The required DFT calculations were carried out with the quantum chemistry software ORCA (version 5.0.3) [94, 95]. We executed spin unrestricted DFT calculations with the PBE exchange-correlation functional [96] in combination with the def2-TZVP basis set [97] and with the $\upomega$B97X functional [98] in combination with the smaller split-valence 6-31G* basis set [99]. The xTB software [100] was applied for semi-empirical GFN2-xTB calculations [101]. For MLP energy, gradient, and Hessian calculations within SCINE, we implemented the module SCINE Parrot (version 1.0.0).
Elementary step trials in the CRN exploration employed the Newton Trajectory Algorithm 2 (NT2) of Chemoton (see Reference [23] for details). In the NT2 scans, uni- and bimolecular reactions were considered for one structure per compound, whereby multiple attack points were enabled, but at most two rotamers and two bond modifications were allowed [87]. We note that structure and compound refer here to their technical terms in SCINE [7]; that is, a structure is a (stationary) point on the potential energy surface and a compound is a group of structures with the same connectivity as defined in SCINE Molassembler. Reactive complexes were not generated for two charged structures with the same sign of the charge. The spin multiplicity of the resulting structures was chosen to be as small as possible. The exploration was only continued for structures which can be reached by an energy barrier up to $250\,\mathrm{kJ\,mol}^{-1}$. $10\%$ of the occurring structure conformations in the NT2 scans as well as in the following intrinsic reaction coordinate (IRC) calculations were stored for the evaluation by the MLPs. In addition, $1\%$ of the structure conformations occurring in the optimizations of reactants, reactive complexes, transition states, and IRC products were saved. The structure conformations were ordered chronologically to enable continual learning retrospectively.
@@ 3.2 Machine Learning Potentials
The eeACSF parameter values of the lMLPs are provided in Tables S1 and S2 in the Supporting Information. Network expressivity by activation rank (NEAR) [106] was applied as a training-free pre-estimator of neural network performance to automate the search for the neural network architecture. The resulting architecture contains 135 input neurons, four hidden layers with 117, 137, 164, and 196 neurons, and one output neuron. The training was based on total energies, from which the sum of the respective reference element energies (Table S3 in the Supporting Information) were subtracted, and on atomic force components. For each lMLP training, we carried out $20$ independent HDNNP training runs, in which the initial neural network weight parameter values and the reference data assignment to training and test sets were different. The weight parameter initialization was tailored to the activation function $\mathrm{sTanh}(x)\coloneqq 1.59223\cdot\tanh(x)$ [47]. In general, $90\%$ of the reference structures were used for training and $10\%$ for testing. Three different schemes for the number of fitted structures per training epoch were applied: (1) a constant fraction of $\tfrac{1}{30}$ of all training structures contained in the respective data set, (2) a constant number of $N_{\mathrm{fit}}=750$ structures, and (3) an adjustment to the number of training structures that are not sorted out (as in Algorithm 1), with a constant fraction $p_{\mathrm{fit}}=\tfrac{1}{30}$. The hyperparameters of lADS are given in Section 2.2. Exceptions were made for scheme (1) $T_{1}=0.75$, $N_{-}=40$, $N_{++}=400$, and $p_{\mathrm{redun}}^{\mathrm{max}}=0.015$, and for scheme (2) $p_{\mathrm{redun}}^{\mathrm{max}}=0.0125$.
In the lMLP weight parameter optimization, a pre-optimization step was carried out based on the energy-dependent term of the loss function $L_{\mathrm{total}}$ (Equation 1) employing each selected structure in a training epoch. Subsequently, the energy and all atomic force components of these structures were utilized in a second optimization step, which included updating the lADS properties. The CoRe optimizer (version 1.1.0) [47, 55, 107] was applied with the hyperparameters $\beta_{1}^{\mathrm{a}}=0.7375$, $\beta_{1}^{\mathrm{b}}=0.8125$, $\beta_{1}^{\mathrm{c}}=250.0$, $\beta_{2}=0.99$, $\epsilon=10^{-8}$, $\eta_{-}=0.55$, $\eta_{+}=1.2$, $s_{\xi}^{0}=10^{-3}$, $s_{\mathrm{min}}=10^{-6}$, $s_{\mathrm{max}}=10^{-2}$, $d=0.1$, $t_{\mathrm{hist}}=250$, and $p_{\mathrm{frozen}}=0.025$. Exceptions were $p_{\mathrm{frozen}}=0$ and $d=0.01$ for the weight parameters $\alpha$ and $\beta$ (see Supporting Information S1.1) and $p_{\mathrm{frozen}}=0$ and $d=0$ for weight parameters associated with the output neuron.
The prediction was based on an ensemble of the $10$ best individual HDNNPs of all $20$. The respective ranking for the selection was given by the sum of the mean squared errors of the test energies and test atomic force components, whereby the energy error was scaled by 2 $2\,500\,\text{\AA}^{-2}$.
@@ 4.1 Chemical Reaction Network
We describe the CRN by technical terms of SCINE [7], where an elementary step is a transformation from a minimum energy structure to another one through a single transition state. A reaction is a group of elementary steps connecting the structures belonging to two reacting compounds (see Section 3.1). Hence, two compounds can be connected by a reaction.
Two CRN explorations were carried out starting both from HCN and H2O. One was based on PBE/def2-TZVP energies and the other on $\upomega$B97X/6-31G* energies. Each exploration proceeded in two shells; that is, all initial elementary step trials (see Section 3.1) were calculated for HCN and H2O, and afterwards, all subsequent trials were calculated for the compounds resulting from the initial trials. The resulting PBE CRN is built from $719$ compounds and 1 23012301\,2301 230 reactions, while the $\upomega$B97X CRN comprises $880$ compounds and 1 50015001\,5001 500 reactions. The difference in these numbers can be explained by the different potential energy surfaces represented by the different exchange-correlation density functionals, which affect the elementary step trials and hence the applied restrictions for the exploration.
@@
The CRN of $\text{HCN}+\text{H}_{2}\text{O}$ is important in the context of the origin of life [108, 109]. The initial two steps in the formation of formic acid (up to the formation of formamide) [110] are contained in the obtained CRN. The last step is expected to be absent because only two exploration shells were considered. Other species such as hydrogen isocyanide, methylene imine, aminoacetonitrile, hydrogen peroxide, and molecular hydrogen are included in the CRN, as are many high-energy species. We note that a proper evaluation of this chemistry would require more exploration shells to fully represent the reaction mechanisms.
@@
However, since we are interested in a typical ensemble of structures that is encountered in CRN explorations, we can limit the exploration to the first two shells. Since the structures that we encounter do not only span minimum-energy and transition-state structures, but also all sorts of positions of the atoms along (not yet optimized) reaction paths, we will call all of them ’structure conformations’ or just ’conformations’ for the sake of brevity in this work. Some reaction coordinates led to structures of comparatively high energies which have to be represented sufficiently well by the MLP. We compiled fractions of the structures occurring in all different exploration calculation steps in a benchmark data set (see Section 3.1). Details on the number of structures from the different calculation steps are provided in Table S4 in the Supporting Information. In total, 225 595225595225\,595225 595 structures were collected from the PBE CRN and 296 093296093296\,093296 093 from the $\upomega$B97X CRN.
4.2 Universal Machine Learning Potentials for Exploration of Chemical Reaction Networks
[…]
$E_{\mathrm{elem}}=\sum_{m=1}^{N_{\mathrm{elem}}}\sum_{n=1}^{N_{\mathrm{atom}}^{m}}\overline{E}_{\mathrm{elem,diff}}^{m}\ ,$ (5)
is taken as a correction in the calculation of the root mean square error (RMSE) of the given MLP energies compared to the reference DFT energies.
[…]
To enable reliable kinetic modeling based on the CRN energies, we require at least chemical accuracy, i.e., $1\,\mathrm{kcal\,mol}^{-1}=4.184\,\mathrm{kJ\,mol}^{-1}$, or better. Since the benchmark data contains systems with up to 12 atoms, a minimum energy accuracy of about $3.614\,\mathrm{meV\,atom}^{-1}$ must be the target. However, the uMLP energies show RMSEs of at least several tens of $\mathrm{meV\,atom}^{-1}$ (Figure 3 (a)), which is an order of magnitude higher than the target. We note that this relative energy difference is also more than an order of magnitude larger than expected if a sufficiently large basis set is replaced by another sufficiently large basis set.
Similarly, the MLP accuracy target of $100\,\mathrm{meV}\,\text{\AA}^{-1}$ for the atomic force components [58] cannot be achieved by the uMLPs for these CRN benchmark data (Figure 3 (b)). Not surprisingly, $\text{HCN}+\text{H}_{2}\text{O}$ CRN explorations based on ANI-2x and M3GNet energies fail because they yield too often unphysical chemical structures and energy profiles. Consequently, the uMLPs investigated are not truly universal and not yet accurate enough for the CRN exploration considered here.

### v2

@@ 2.2 Lifelong Adaptive Data Selection
If the exclusion strike counter $X^{r}$ reaches its limit $N_{\mathbf{X}}$, $S_{\mathrm{hist}}^{r,E/F}$ is set to 3-3 and the structure is excluded due to very large errors. If the upper threshold for $S_{\mathrm{hist}}^{r}$ as defined in Algorithm 2 is exceeded, 2-2 is assigned to $S_{\mathrm{hist}}^{r,E/F}$ to exclude the structure due to steadily large errors for many training steps. Both assignments mean that the respective training data point is inconsistent with the majority of data. If the lower threshold for $S_{\mathrm{hist}}^{r}$ is exceeded, $S_{\mathrm{hist}}^{r,E/F}$ is set to 1-1, classifying a redundant structure, i.e., a structure that has been well represented for many steps. To avoid that too many structures are classified as redundant in the same training step, a maximal fraction of the new redundant training structure per step $p_{\mathrm{redun}}^{\mathrm{max}}$ can be set in addition to Algorithm 2. In this way, only
@@
randomly selected structures of the structures exceeding the lower threshold are classified as redundant and $S_{\mathrm{hist}}^{r,E/F}$ of the remaining structures is divided by $F_{--}$. This approach reduces the risk that all structures creating some redundancy are by chance excluded in the same step, i.e., none of them remains available for training.
To update the fraction of good data $p_{\mathrm{good}}$, the new loss contributions first replace the respective old ones in $\mathbf{L}_{\mathrm{old}}^{E/F}$. Subsequently, the mean $\overline{L}_{\mathrm{new}}^{E/F}$ of this vector is calculated for the union of available and redundant training data. If the new means of energies and forces are not larger than the respective old values, $p_{\mathrm{good}}$ will decrease by $p_{\mathrm{good}}^{\mathrm{max}}\,N_{p}^{-1}$. Otherwise, it increases by the same value. The mean loss contributions for energy and forces are initialized before the first step as infinity. The resulting $p_{\mathrm{good}}$ has a lower bound of 0 and an upper bound of $p_{\mathrm{good}}^{\mathrm{max}}$. Consequently, $p_{\mathrm{good}}$ can have $N_{p}+1$ different values. Finally, $\overline{L}_{\mathrm{old}}^{E/F}$ is overwritten by $\overline{L}_{\mathrm{new}}^{E/F}$.
@@
For an efficient integration of additional data at a training stage where some structures are already well represented, Algorithm 3 can replace lines 3 to 6 of Algorithm 2. The main idea behind this data integration is that the additional data can have the maximum training probability for several training steps and they cannot be excluded during these steps. In this way, the lMLP can learn the new structures fast, even if the representation quality is very different between the previous and new training data. Without this integration algorithm, the risk of exclusion is high for new data because the typically low errors for the majority of old data lead to a fast increase in the exclusion strike counters and the adaptive selection factors of new data with usually high errors. In addition, the new data initially do not affect the update of the selection-determining properties of the old data to obtain a stable assessment of the representation quality.
@@
[…]
@@
$\mathbf{L}_{\mathrm{new}}^{E/F}\leftarrow\mathrm{NaN}\ \ \mathrm{if}\ 0<I^{r}<N_{\mathbf{I}}$
@@
$\mathbf{t}_{\mathrm{BT}}\leftarrow\mathrm{cumsum}\left\{\left[\mathrm{min}\left(\dfrac{N_{\mathrm{fit}}\cdot\mathbf{P}^{r^{\prime}}}{P^{\mathrm{sum}}+\mathbf{P}^{r^{\prime}}},\ 1\right)\right]^{-1}\right\}$
@@
The following part of Algorithm 4 has to be repeated for each structure $r^{\prime}$ that has been identified in the current step to be inconsistent. Initially, the non-normalized probabilities $\mathbf{P}^{r^{\prime}}$ are calculated using the current $P_{\mathrm{loss}}^{r^{\prime}}$, $p_{\mathrm{good}}$, and $\mathbf{F}_{\mathbf{S}}$ or $\mathbf{F}_{\mathbf{X}}$. The latter depends on which exclusion reason applies for $r^{\prime}$. Afterwards, the average training frequency is calculated for each entry in $\mathbf{P}^{r^{\prime}}$, whereby a structure can only be trained once in a training step. The cumulative sum minus the value of the first vector entry yields the vector of backtracking steps $\mathbf{t}_{\mathrm{BT}}$. The current epoch $n_{\mathrm{epoch}}$ minus $\mathbf{t}_{\mathrm{BT}}$ corresponds to the most probable training steps of structure $r^{\prime}$ under the aforementioned assumptions. For these steps, the decay hyperparameters $\boldsymbol{\upbeta}_{1}$ of the CoRe optimizer are calculated. Subsequently, the backtracking factors $F_{\mathrm{BT}}^{1}$ and $F_{\mathrm{BT}}^{2}$ can be obtained, i.e., the sum of the weights of all loss gradient contributions of structure $r^{\prime}$ in the steps $n_{\mathrm{epoch}}-\mathbf{t}_{\mathrm{BT}}$ can be calculated using the decay hyperparameters $\boldsymbol{\upbeta}_{1}$ and $\beta_{2}$.
@@
For the lADS hyperparameters, we generally recommend the following settings: $p_{\mathrm{fit}}=\tfrac{1}{30}$, $S_{\mathrm{hist}}^{\mathrm{min}}=0.1$, $S_{\mathrm{hist}}^{\mathrm{max}}=100$, $T_{i}=\{0.9,2.5,4.0,7.5\}$, $N_{--}=10$, $N_{-}=30$, $N_{+}=100$, $N_{++}=300$, $N_{\mathbf{X}}=15$, $p_{\mathrm{redun}}^{\mathrm{max}}=0.02$, $p_{\mathrm{good}}^{\mathrm{max}}=\tfrac{2}{3}$, $N_{p}=20$, and $N_{\mathbf{I}}=30$. The data integration algorithm (Algorithm 3) can be applied as soon as new data are added for the first time. Hence, Algorithm 3 needs to be disabled only if training is started from scratch. In addition, the hyperparameter for balancing the energy and force loss contributions was $q=250$ in this work. These settings were applied unless otherwise stated.
@@ 3.1 Exploration of Chemical Reaction Networks
For the CRN exploration we applied our freely available open-source Software for Chemical Interaction Networks (SCINE) (see Reference [79] for an overview), especially the SCINE modules Chemoton (version 3.0.0) [23, 80], Molassembler (version 2.0.0) [81, 82], ReaDuct (version 5.0.0) [83, 84], Puffin (version 1.2.0) [85], and Database (version 1.2.0) [86]. The required DFT calculations were carried out with the quantum chemistry software ORCA (version 5.0.3) [87, 88]. We executed spin unrestricted DFT calculations with the PBE exchange-correlation functional [89] in combination with the def2-TZVP basis set [90]. The xTB software [91] was applied for semi-empirical GFN2-xTB calculations [92]. For MLP energy, gradient, and Hessian calculations within SCINE, we implemented the module Parrot (version 1.0.0).
To describe the setup of the CRN exploration, we refer here to technical terms in SCINE [7]: A structure is a (stationary) point on the potential energy surface and a compound is a group of structures with the same connectivity as defined in Molassembler. An elementary step is a transformation from a minimum energy structure to another one through a single transition state. A reaction is a group of elementary steps connecting the structures belonging to two reacting compounds. Hence, two compounds can be connected by a reaction.
Elementary step trials in the CRN exploration were set up for one structure per compound, i.e., we did not explore explicitly changes in the reactivity due to different conformations of a compound. However, for bimolecular reaction trials, the required reactive complexes were generated using different attack points of the two structures and up to two rotamers of them. Reactive complexes were ignored if the two structures were both charged positively or negatively, since electrostatic repulsion makes the resulting elementary steps unfavorable. To find the transition state, we employed the Newton Trajectory Algorithm 2 (NT2) (see Reference [23] for details). We allowed uni- and bimolecular reaction trials with one or two intended bond modifications per trial. We note that the resulting number of bond modifications can be lower or higher, since the NT2 scan is a single-ended transition state search algorithm. If an NT2 scan found a transition state, the respective structure was optimized. Afterwards, an intrinsic reaction coordinate (IRC) calculation was carried out to get the reactant and product minimum energy structures that belong to the transition state. The spin multiplicity of the resulting structures was chosen to be as small as possible. Structures that can only be formed by overcoming a single energy barrier larger than $250\,\mathrm{kJ\,mol}^{-1}$ were not considered in the setup of further elementary step trials. $10\%$ of the structure conformations that occurred in the NT2 scans and IRC calculations were stored for the benchmark of the lMLPs. In addition, $1\%$ of the structure conformations occurring in the optimizations of reactants, reactive complexes, transition states, and IRC outcomes were saved. The structure conformations were ordered chronologically to enable continual learning retrospectively.
@@ 3.2 Machine Learning Potentials
We implemented training and prediction of lMLPs in the lMLP software (version 2.0.0). It utilized NumPy (version 1.26.4) [93], PyTorch (version 2.3.1) [94], and Numba (version 0.60.0) [95].
The eeACSF parameter values of the lMLPs are provided in Tables S1 and S2 in the Supporting Information. Network expressivity by activation rank (NEAR) [70] was applied as a training-free pre-estimator of neural network performance to automate the search for the neural network architecture. The resulting architecture contains 135 input neurons, four hidden layers with 117, 137, 164, and 196 neurons, and one output neuron. The training was based on total energies and the atomic force components. The sum of respective reference element energies (Tables S3 and S4 in the Supporting Information) was subtracted from the total energy prior to training. In this work, we trained on PBE/def2-TZVP energies and forces, but, in general, data from any electronic structure method (including dispersion corrections) can be utilized, provided that the potential energy surface is consistent. For each lMLP training, we carried out $20$ independent HDNNP training runs, in which the initial neural network weight parameter values and the reference data assignment to training and test sets were different. The weight parameter initialization was tailored to the activation function $\mathrm{sTanh}(x)\coloneqq 1.59223\cdot\tanh(x)$ [58]. In general, $90\%$ of the reference structures were used for training and $10\%$ for testing.
In the optimization of the lMLP weight parameters, a pre-optimization step was carried out based on the energy-dependent term of the loss function $L_{\mathrm{total}}$ (Equation 1), since this term is much faster to evaluate than the force-dependent term. The pre-optimization step employed the energies of all structures that were selected by lADS for the given training epoch. Subsequently, the energies and atomic force components of these structures were utilized in a second optimization step, which included updating the lADS properties. The hyperparameters of lADS matched our recommendations in Section 2.2. To obtain fair comparisons between continual and conventional learning, exceptions were made in Sections 4.3 and 4.4 that are explicitly stated there. The CoRe optimizer (version 1.1.0) [58, 66, 96] was applied with the hyperparameters $\beta_{1}^{\mathrm{a}}=0.7375$, $\beta_{1}^{\mathrm{b}}=0.8125$, $\beta_{1}^{\mathrm{c}}=250.0$, $\beta_{2}=0.99$, $\epsilon=10^{-8}$, $\eta_{-}=0.55$, $\eta_{+}=1.2$, $s_{\xi}^{0}=10^{-3}$, $s_{\mathrm{min}}=10^{-6}$, $s_{\mathrm{max}}=10^{-2}$, $d=0.1$, $t_{\mathrm{hist}}=250$, and $p_{\mathrm{frozen}}=0.025$. Exceptions were $d=0.01$ for the weight parameters $\alpha$ and $\beta$ (see Supporting Information S1.1 for the definition of these weight parameters) and $d=0$ for weight parameters associated with the output neuron, since these weight parameters can intentionally have larger values. Additionally, we set $p_{\mathrm{frozen}}=0$ for these weight parameters.
@@ 4.1 Chemical Reaction Network
A CRN exploration was carried out starting from HCN and H2O. The exploration proceeded in two shells; that is, all initial elementary step trials (see Section 3.1 for the definitions of the technical terms of SCINE [7]) were calculated for HCN and H2O, and afterwards, all subsequent trials were calculated for the compounds resulting from the initial trials. The resulting CRN is built from $719$ compounds and 1 2301\,230 reactions.
@@
The CRN of $\text{HCN}+\text{H}_{2}\text{O}$ is important in the context of the origin of life [97, 98]. The initial two steps in the formation of formic acid (up to the formation of formamide) [99] are contained in the computed CRN. The last step is expected to be absent because only two exploration shells were considered. Other species such as hydrogen isocyanide, methylene imine, aminoacetonitrile, hydrogen peroxide, and molecular hydrogen are included in the CRN, as are many high-energy species. We note that a proper evaluation of this chemistry would require more exploration shells to fully represent the reaction mechanisms.
@@
However, in this work we are mainly interested in a typical ensemble of structures that is encountered in CRN explorations. Therefore, we can limit the exploration to the first two shells. Since the structures that we encounter do not only span minimum-energy and transition-state structures, but also all sorts of positions of the atoms along (not yet optimized) reaction paths, we will call all of them ‘structure conformations’ for the sake of brevity in this work. Some reaction coordinates led to structures of comparatively high energies which have to be represented sufficiently well by the MLP. We compiled fractions of the structure conformations occurring in the elementary step trials in a benchmark data set. Details on the number of structure conformations obtained from the different subtasks of the elementary step trials are provided in Table S5 in the Supporting Information. In total, 225 595225\,595 structure conformations were collected from the CRN exploration.